x^2+3x-40=0
How would you solve this equation?

Answer by kapilsinghi(68)

x^2+3x-40=0
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x-5)(x+8)=0
x=5 or x=-8
kapilsinghi123@gmail.com